OpenStudy (ashy98):

2 added to x is 4. Find x.

OpenStudy (anonymous):

2 + x = 4 x = 2

OpenStudy (anonymous):

2 added to x = x+2 x+2 = 4 x=2

OpenStudy (ashy98):

thanks (: